1.Clydesdale Bank is setting up a brand new branch. The cost of the project will be $1.2 million. The branch will create additional cash flows of $235,000, $412,300, $665,000 and $875,000 over the next four years. The company's cost of capital is 12 percent. What is the internal rate of return on this branch expansion? (Round to the nearest percent.) HINT: Use a financial calculator or the Excel function: =IRR(values,[guess])

Select one:

A. 20%

B. 23%

C. 25%

D. 27%

2.A fast growth share has the first dividend (t=1) of $3.28. Dividends are then expected to grow at a rate of 9 percent p.a. for a further 4 years. It then will settle to a constant-growth rate of 1.9 percent. . If the required rate of return is 16 percent, what is the current price of the share? (to the nearest cent)

Select one:

a. $23.26

b. $25.88

c. $28.46

d. $83.72

3.An active secondary market enables companies to sell their new debt or equity issues at

Select one:

A. higher funding costs

B. equal funding costs

C. lower funding costs

D. none of the above
Answer #1

1.

The IRR can be calculated as mentioned below:

2 Time 30 4 1 52 63 7 4 8 IRR Cashflow - 1200000 235000 412300 665000 875000 =IRR(B3:B7,5%)

ДА В 2 Time 3 4 5 Cashflow 0-1200000 1 235000 2 412300 3 665000 4 875000 23% 6 7 8 IRR

The IRR is 23%.

The correct option is B.

2.

1 Required return 2 Time 3 1 42 5 3 64 75 0.16 Dividend 3.28 =B3*(1+$D$1) =B4*(1+$D$1) =B5*(1+$D$1) =B6*(1+$D$1) Growth rate

B D 9% 1 Required return 2 Time u wn 16% Growth rate Dividend Discount factor Present value 3.28 0.862068966 2.83 3.58 0.7431

The current price is $28.46.

The correct option is c.

3.

Company can sell their new debts or equities at lower cost.

The correct option is C.